A waterbed heater uses 425W of power. It is on 35% of the time, off 65% . What is the annual cost of electricity at a billing rate of$0.13 per kWhr?
Answer #1

1 year=8765.81 hrs=24*365

Out of which only 35%,it is on

so on time=3068.03 hrs

Wattatge=0.425 kw

The kwH=0.425*3068.03=1303.914

Cost=1303.914*0.13=169.50 $
Answer #2

No. of hours in a year =8765.81

its 35% = 35/100*8765.81

= 3068.03 hr

425 W in an hour

therefore for one year it will use =3068.03 hr * 425 W

= 1303.914 KWhr

therfor total coast = $ 0.13 per KWhr * 1303.914 KWhr

= $ 169.508